rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#5198 — FS#9122 — Improvement Lille 10G optic loop

Attached to Project— ADSL / SDSL
Modernization
Backend / Core
CLOSED
100%
We are going to insert a new Lambda into the fibre optic to increase the capacity on our NRAs.
We are working on this loop: PIL59 - BLE59 - BOI59 - FAC59 - SEC59 - LIB59.
No impact on traffic is expected. Works will commence Friday August 9th at 14:00.
Date:  Tuesday, 20 August 2013, 15:51PM
Reason for closing:  Done
Comment by OVH - Friday, 09 August 2013, 15:54PM

We are starting the maintenance works, we are inserting the 10G cards on the relevant sites.


Comment by OVH - Friday, 09 August 2013, 15:55PM

We are working on the first side of the loop.


Comment by OVH - Friday, 09 August 2013, 18:47PM

We have an optic attenuation issue. We are halting the intervention. Optimisation works are required on the optic loop. We will start again on Monday 12th.

Securisation is operational in case of power failure.


Comment by OVH - Monday, 12 August 2013, 16:23PM

We are continuing with the intervention on the Lille optic loop. We
are going to reconfigure the EDFA amps in order to move the loop
onto 10G.


Comment by OVH - Monday, 12 August 2013, 17:12PM

We are moving the amp to boi59, securisation is operational in case of power failure.


Comment by OVH - Friday, 16 August 2013, 13:07PM

We have succeeded to raise one side of the 10G loop,
however, we are still having reception issues with securisation.
We are going to intervene this afternoon in order to
perform tests on the fibre optic.
We will start the intervention at 3pm Friday August 16th.


Comment by OVH - Friday, 16 August 2013, 18:08PM

We have found the issue on the securisation side, the
optic levels were too weak. We will intervene on Monday
August 19th to insert an EDFA amp, in order to obtain
the correct optic power.